What is the probability that the shelf life of the




battery is between 500 and 550 days?





1
Expert's answer
2022-05-03T03:54:03-0400

A certain type of battery has a mean shelf life of 600 days with a standard deviation of 28 days: "X\sim N(600,28^2)."


"P(500<X<550)=P(Z<\dfrac{550-600}{28})"

"-P(Z\le\dfrac{500-600}{28})"

"\approx P(Z<-1.7857)-P(Z\le-3.5714)"

"\approx0.03707-0.00018"

"\approx0.0369"